OpenStudy (anonymous):

Simplify the following expressions and write the results in the polar form re iθ . ( √ 3 + i ) ^5

mathslover (mathslover):

can you write it by using equation facility?

OpenStudy (anonymous):

your job is to write \(\sqrt{3}+i\) in trig (polar) form do you know how to do that? you need two numbers, \(r\) and \(\theta\) to write \(a+bi\) as \(re^{i\theta}\) \(r=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

in your case \(r=\sqrt{\sqrt{3}^2+1^2}=\sqrt{3+1}=\sqrt{4}=2\)

OpenStudy (anonymous):

as for \(\theta\) you can sometimes use \(\theta=\tan^{-1}(\frac{b}{a})\) if you are in quadrant 1 or 4 in your case you get \(\theta=\tan^{-1}(\frac{1}{\sqrt{3}})\)

OpenStudy (anonymous):

do you know that number?